OpenStudy (anonymous):

how do i find the inverse of the function f(x)=5 . log(2x + 3)

OpenStudy (mathmate):

switch x and y and solve for y in terms of the new x. x=5log(2y+3) log(2y+3)=(x/5) 2y+3=e^(x/5) y=(e^(x/5)-3)/2 Check that f-1(f(x))=x